### v1

# Extreme Jet Beaming Observed in Neutrino-Associated Blazars

Abstract
Bright blazars were found to be prominent neutrino sources, and a number of IceCube events were associated with them over recent years. A particularly strong observational connection is present between neutrinos and blazars with bright, Doppler-boosted, parsec-scale radio emission. In this work, we further explore the nature of this connection by examining the jet geometry and kinematics of neutrino-associated blazars. We find that these blazars demonstrate remarkably strong jet beaming, even compared to other radio-bright sources. Their Doppler and Lorentz factors are larger, and viewing angles are smaller than for other blazars in the complete uniformly selected MOJAVE sample. Observationally, this serves as yet another piece of evidence for blazars forming a major population of neutrino sources. The strong neutrino-beaming correlation indicates that high-energy neutrino velocity is predominantly oriented along the jet, and the original PeV-scale protons exhibit a relativistic bulk motion along the jet. It suggests that neutrino production happens not too close to the black hole, but rather at sub-parsec distances, where the jet is already accelerated.
journal: ApJ
1 Introduction
Blazars were the first high-energy (TeVs-PeVs) astrophysical neutrino sources to be discovered (IceCube Collaboration et al., 2018). Later, neutrinos have also been detected from other sources, including the Milky Way Galaxy (Kovalev et al., 2022; Icecube Collaboration et al., 2023; Albert et al., 2023; Allakhverdyan et al., 2024a), nearby Seyfert-type Active Galactic Nuclei (AGNs, IceCube Collaboration et al., 2022; Neronov et al., 2024), and Tidal Disruption Events in AGNs (TDEs, van Velzen et al., 2024). Still, blazars remain a very prominent class of bright neutrino emitters. There are more and more population-based studies resulting in neutrinos associated with AGNs selected by their compact radio flux (basically, blazars; e.g., P23; Kouch et al. 2024; Albert et al. 2024; Suray & Troitsky 2024), X-ray flux (Kun et al. 2024; Khatee Zathul et al. 2024; Neronov et al. 2024; blazars in Plavin et al. 2024), or using more specialized blazar selection methods (e.g., Kadler et al., 2016; Giommi et al., 2020; Buson et al., 2022).
Recent studies (e.g. P23; Kouch et al. 2024) have specifically demonstrated a strong observational connection of radio-bright blazars being also bright neutrino sources. This connection is manifested both as the neutrino – pc-scale radio flux correlation (see Section 3 for updates), and also as a temporal correlation with neutrinos coming predominantly during radio flares. Those results already provide an indication that neutrinos are emitted predominantly along the jet axis in blazars; more direct studies should help quantifying this effect. Furthermore, note that the strong neutrino-blazars association is not found in every analysis, depending on the handling of directional uncertainties (see discussion in Abbasi et al., 2023; Kouch et al., 2024). Utilizing additional observables is thus crucial to refine and potentially confirm the connection.
With the growing sample of neutrino-coincident blazars, we can investigate the nature of the jet-related effects on the neutrino emission in more detail. In this paper, we focus on a direct study of the jet geometry and kinematics of blazars that are likely neutrino sources. Specifically, we explore beaming indicators, such as the Doppler factor and the jet viewing angle, for blazars that coincide with high-energy neutrinos detected at the IceCube (Section 2). This comparison enables us to establish more directly that neutrinos are effectively beamed, similar to electromagnetic emission (Section 4.1). We discuss the implications of these findings for constraining the neutrino production region and the jet acceleration mechanisms (Section 4.2, Section 5). We demonstrate the robustness of the findings presented in this paper and of our earlier results by the consistency between different IceCube event samples in Section 3 and Section 4.1.
2 Data
Figure 1: The sky distribution of radio blazars and IceCube events used in this work. See Section 2 for the definitions of each sample. Two different IceCube event selections are shown in different manners (filled and unfilled) to enable a clearer visual comparison between them. There are 83 blazars from the RFC catalog coincident with at least one IceCube neutrino; note that they include both real physical associations and chance coincidences, see Figure 2. Neutrino-coincident MOJAVE 1.5JyQC (Section 2.1) blazars are labeled here; see Figure 3 for their VLBI images and Table 1 for their properties.
2.1 Radio observations of parsec-scale emission
[…]
For Doppler boosting and jet viewing angle measurements, we rely on the results of the MOJAVE (Monitoring Of Jets in Active galactic nuclei with VLBA Experiments) team. These parameters cannot be measured directly for blazars; the MOJAVE team has suggested (Homan et al., 2006) and applied (Homan et al., 2021) a procedure to infer them from VLBI monitoring data; we use those measurements in this paper. The full MOJAVE dataset is assembled with a range of selection criteria: bright VLBI-selected blazars, compact symmetric objects, Fermi-selected blazars, low-luminosity AGNs (Lister et al., 2011, 2019, 2020, 2013). Similarly to our earlier studies, we focus on a uniform flux-density limited subsample of the MOJAVE program sources in the analysis: the 1.5 Jy Quarter Century Sample (1.5JyQC; defined in Lister et al. 2019). It consists of blazars at $\delta>-30\degree$ with VLBI flux density exceeding 1.5 Jy at 15 GHz at least once during the quarter century of observations, 1994–2019. This sample contains 232 blazars, and 205 of those have Doppler boosting measurements (highlighted in Figure 1). The MOJAVE 1.5JyQC sample was specifically designed to be homogeneous and unbiased.
2.2 IceCube high-energy neutrinos
From the neutrino side, we utilize the high-energy events reported by the IceCube observatory. We consider two high-energy event catalogs, both to ensure the consistency of our results and keep the continuity with earlier studies. Both datasets are analyzed in the exact same way, and we ensure that all reported results are consistent between the two.
The P23 IceCube event selection.
The first catalog is the sample of high-energy alert-like neutrino detections we collected before, most recently in P23. Earlier, the same selection criteria were utilized in P20 and Hovatta et al. (2021). This sample consists of 71 events with the angular uncertainty region below 10 sq. deg. and the estimated energy above 200 TeV.
The IceCat-1 event catalog.
In 2023, the IceCube collaboration released the IceCat-1 track event catalog (IceCube Collaboration, 2023, last updated in November 2023). It is mostly based on a different reconstruction of IceCube detection events, utilizes different event selection and categorization criteria. Following the IceCube guidance on event categories in the catalog, we select 93 Gold events from the IceCat-1 with the angular uncertainty region below 10 sq. deg. The estimated signalness is $\geq 50\%$ for all these events; their neutrino energy estimates range from 56 TeV to 6 PeV.
For the analyses performed in this work, we enlarge the original IceCube directional 90% uncertainties by $0.78\degree$ isotropically in all directions. This enlargement was determined in P23 to be optimal for association searches, and we keep the same value to avoid multiple trials in the statistical analysis. A similar enlargement was found in Kouch et al. (2024); see that paper, P23, and Section 3 for a more detailed discussion.
The sky localization of IceCube events from both samples is shown in Figure 1. Blazars coincident with IceCube events are highlighted in the figure, and those from the MOJAVE 1.5JyQC sample are labeled; see Table 1 for their properties.
Figure 2: The distribution of VLBI 8 GHz flux density for blazars coincident with IceCube neutrinos, shown as histograms and cumulative distribution plots. Similarly to Figure 1, all VLBI blazars from the RFC catalog are shown in black exactly the same in the left and right panels. Neutrino coincidences are shown in red, using the P23 events sample on the left and the IceCat-1 on the right (see Section 2.2).
3 Neutrino-blazar correlation
Table 1: Key properties of high-energy neutrino coincident blazars that are observed within the MOJAVE program.
Neutrino event Blazar name VLBI Flux Beaming Indicators
Date Event ID J2000 B1950 $z$ 8 GHz 15 GHz 1.5JyQC $D$ $\Gamma$ $\theta$
(P23 sample) (IceCat Gold sample) (Jy) (Jy) ()
(1) (2) (3) (4) (5) (6) (7) (8) (9) (10) (11)
2015--09--26 IC150926A J1256--0547 1253--055 $0.536$ $15$ $19$ yes $140$ $72$ $0.12$
2022--03--03 J1751$+$0939 1749$+$096 $0.322$ $2.8$ $4$ yes $100$ $51$ $0.07$
2019--07--30 IC190730A J1504$+$1029 1502$+$106 $1.838$ $1.5$ $1.4$ yes $65$ $35$ $0.44$
2016--01--28 IC160128A J1733--1304 1730--130 $0.902$ $4$ $4.5$ yes $48$ $32$ $1$
2015--08--12 IC150812B J2148$+$0657 2145$+$067 $0.999$ $6.6$ $7.5$ yes $31$ $16$ $0.36$
2015--08--31 IC150831A J0336$+$3218 0333$+$321 $1.259$ $1.6$ $1.7$ yes $30$ $18$ $1.4$
2012--05--15 IC120515A J1310$+$3220 1308$+$326 $0.997$ $1.9$ $2.1$ yes $24$ $28$ $2.3$
2022--02--05 IC220205B J1743--0350 1741--038 $1.054$ $4$ $5.3$ yes $22$
2011--09--30 J1743--0350 1741--038 $1.054$ $4$ $5.3$ yes $22$
2013--10--14 IC131014A J0211$+$1051 0208$+$106 $0.2$ $0.66$ $1.1$ yes $6.3$ $5.1$ $8.8$
2020--11--30 IC201130A J0206--1150 0203--120 $1.663$ $0.28$ $0.3$ no $35$ $20$ $1$
2017--09--22 IC170922A J0505$+$0459 0502$+$049 $0.954$ $0.7$ $0.93$ no $25$ $14$ $1.6$
IC221223A J2311$+$3425 2308$+$341 $1.817$ $0.69$ $0.86$ no $23$
IC201221A J1724$+$4004 1722$+$401 $1.049$ $0.5$ $0.58$ no $9.8$ $61$ $3.2$
2010--10--09 J2200$+$1030 2157$+$102 $0.172$ $0.19$ $0.22$ no $5$
2017--09--22 IC170922A J0509$+$0541 0506$+$056 $0.3365$ $0.42$ $0.46$ no $1.8$ $1.5$ $33$
2011--07--14 IC110714A J0432$+$4138 0429$+$415 $1.022$ $1.4$ $1$ no $0.13$ $54$ $31$
Note. — Columns are as follows: (1) — IceCube neutrino arrival date for events included in the P23 sample; (2) — IceCat event ID for events included in IceCat-1; see Section 2.2 for a definition of these samples; (3), (4) — coordinate-based blazar names, with the J2000 name taken directly from the RFC (Petrov & Kovalev, 2025); (5) — redshift $z$ extracted from the NASA/IPAC Extragalactic Database (NED); (6), (7) — parsec-scale radio flux densities measured by VLBI, as presented in the RFC catalog; their typical uncertainty is about $10\%$, dominated by the amplitude calibration; (8) — flag indicating whether the blazar is included in the uniformly-selected MOJAVE 1.5JyQC sample (Section 2.1); those sources are also labeled in Figure 1, and their VLBI images are shown in Figure 3; (9), (10), (11) — beaming indicators, Doppler $D$ and Lorentz $\Gamma$ factors and the viewing angle $\theta$, as measured within the MOJAVE program (Homan et al., 2021); see Section 4.1 for a discussion of assumptions and uncertainties of these measurements.
The blazars are ordered by the inclusion in MOJAVE 1.5JyQC (8) and by the Doppler factor measurement (9).
The statistical correlation between VLBI-bright blazars and high-energy neutrinos has already been established and, as confirmed in our previous work, strengthens over time with more neutrino detections (most recently, P23). In this work, we do not aim to repeat or reevaluate that analysis. Instead, we utilize another dataset of IceCube events (IceCat-1, Section 2.2) to serve as an additional check of the robustness of our findings. Both the IceCat-1 dataset and the neutrino events sample from P23 are used in the following sections of this paper.
[…]
After having chosen the test statistic, we calculate the probability that its observed value could have arisen by chance, assuming no directional neutrino-blazar correlation. For that, we randomize the Right Ascensions of all IceCube events $10^{6}$ times, motivated by the instrument sensitivity being dependent on the Declination only (used in Aartsen et al., 2017, and elsewhere). Then, the chance probability is the fraction of random realizations where the average VLBI flux of matching blazars is higher than the observed value. There are no free parameters to be optimized in this procedure, and we can immediately interpret the resulting fraction as the $p$-value.
The P23 IceCube event selection.
The VLBI blazar-neutrino correlation for this event sample was performed in P23, resulting in a $3.6\sigma$ significance ($4.2\sigma$ pre-trial). For ease of comparison and to remind the reader, we visualize this result in Figure 2.
The IceCat-1 event catalog.
To validate our earlier findings, we also utilize the IceCat-1 Gold dataset released by IceCube in 2023 (Section 2.2). Applying the same statistical procedure, we obtain the chance probability of $p=2.1\%$ ($2.4\sigma$; no trials) as shown in Figure 2. The correlation remains convincing, albeit with a significance level lower than for our previous event selection. We find such an agreement between results obtained from different event reconstructions, selections, and categorization approaches quite remarkable.
Note that updated or more detailed event reconstructions are not necessarily better in every aspect: see demonstrations and discussion in the review Troitsky (2024) and in Sommani et al. (2024). For example, a reconstruction update lead to a decrease of the TXS 0506+056 flare significance (IceCube Collaboration et al., 2021; Abbasi et al., 2023). A detailed analysis and interpretation of the differences between various IceCube event reconstructions is out of scope of this work.
Overall, we see evidence for neutrino-blazar correlation using either the P23 or the IceCat-1 samples. Both give consistent results and are used in the following sections of this paper for more detailed studies. This helps us ensure that we only report findings that hold across different neutrino event samples.
Cross-check without enlarging IceCube errors.
As detailed in Section 2.2, we enlarge original IceCube event uncertainties by $0.78\degree$ for all calculations in this paper, following the approach of P23 and Kouch et al. (2024) and avoiding multiple comparisons issues. Here, we present a focused cross-check that such an enlargement remains important for both IceCube event catalogs we use (Section 2.2).
Repeating our statistical testing procedure with original IceCube uncertainties taken as-is, we don’t see any significant evidence for the correlation with VLBI blazars. The chance probability $p>5\%$ for both IceCube event selections. The spatial correlation shown above is heavily driven by blazars that lie slightly outside the original reported 90% directional errors. This is still consistent with the observed correlation caused by a physical connection. Such an effect could result from unaccounted systematic uncertainties or other discrepancies between estimated and actual error regions. A detailed dedicated study can only be performed by the IceCube collaboration, utilizing all the available data about each detection; see Lagunas Gualda et al. (2021) for a discussion and an overview of challenges. Further in the future, it will be crucial to compare the brightest neutrino sources in the sky between different neutrino observatories. The cross-check we performed here does not replace the detailed analysis performed in P23 and Kouch et al. (2024), but remains consistent with those works and results reported by the IceCube (Abbasi et al., 2023, 2024a).
4 Geometry and kinematics of neutrino-associated blazars
Figure 3: MOJAVE VLBI images for the nine 1.5JyQC blazars coincident with IceCube neutrinos from both samples of Section 2. We show here averages of multi-decade individual VLBI images, which provide a better dynamic range to trace the overall structure of the jet. These stacked images are available from the MOJAVE website (https://www.cv.nrao.edu/MOJAVE/), see details in Pushkarev et al. (2017). For each blazar, we show the redshift $z$, the Doppler factor $D$ and the viewing angle $\theta$ estimates. Blazars are ordered by their Doppler factor, same as in Table 1.
Figure 4: Comparison of observational jet beaming indicators (Doppler $D$ and Lorentz $\Gamma$ factors, viewing angle $\theta$) for neutrino-coincident blazars and for the complete MOJAVE 1.5JyQC sample (see Section 2 and Table 1). Shown are cumulative distributions for all three parameters (the top row). The Doppler factor is among the most direct observational tracers, and we show its histogram with individual neutrino coincidences marked below. Note: for the P23 neutrino sample, there are 9 coincident blazars; one of them is coincident with two neutrinos, adding up to 10 coincidences. The chance probabilities to obtain the observed Doppler factor excess for neutrino coincidences are shown in the plot (see Section 4 for details).
4.1 Observational evidence for neutrino beaming
The association of high-energy neutrinos with VLBI-bright blazars (Section 3) already suggests that beaming plays a major role in neutrino production and detection: neutrinos are predominantly emitted along the direction of the jet. Indeed, VLBI-selected AGNs are dominated by galaxies with jets pointing towards us, their typical viewing angles are $\theta=2\dots 3\degree$ (Lister et al., 2019; Homan et al., 2021). If neutrino emission were isotropic, we would expect to see thousands ($30004\pi/\pi\theta^{2}~{}\sim~{}$ 3000) of similar AGNs pointed in other directions for each neutrino-VLBI association. This is clearly incompatible with the observed number of high-energy neutrinos.
To further probe and quantify the beaming of neutrino emission more directly, we explore kinematic and geometric properties, measured for individual blazars. Specifically, we analyze the plasma Lorentz factor $\Gamma$ (the bulk jet velocity), the viewing angle $\theta$ (between the jet direction and the Earth), and the Doppler factor $D$ (a key measure of relativistic beaming). These three quantities characterize the plasma flow geometry in different but related ways:
$\begin{split}\Gamma&=\frac{1}{\sqrt{1-\beta^{2}}}\\ D&=\frac{1}{\Gamma(1-\beta\cos\theta)}\end{split}$ (1)
They are measured from VLBI observations within the MOJAVE program (Section 2.1, Homan et al. 2021). These estimates crucially rely on the VLBI ability to resolve the parsec-scale structure and determine the brightness of the most compact detail – the so-called core. Due to the measurement procedure, the estimates of these quantities are somewhat related to the VLBI flux density we use in the correlation analysis above (Section 3). Still, they are not exactly the same and differ in fundamental ways. The VLBI flux density is a measure of the total radio emission on multi-parsec scale, while Doppler factors are determined from the brightness of even more compact regions, typically a fraction of a parsec in projection. Viewing angle and Lorentz factor estimates additionally rely on parsec-scale kinematics to constrain the geometry. It can be challenging to calculate precise uncertainties of these measurements. Still, as a cross-check, Doppler factors resulting from MOJAVE measurements were found to be consistent with independent estimates obtained via different techniques, such as radio variability (Liodakis et al., 2018; Homan et al., 2021).
We probe the jet beaming — neutrino emission connection by comparing these beaming indicators (Doppler factor, plasma velocity, viewing angle) between neutrino-coincident blazars and the overall sample. To ensure maximal robustness of results, we remain conservative and focus on a uniform flux density-limited MOJAVE subsample, 1.5JyQC (Section 2.1). The comparison of blazars close to neutrino arrival directions with the whole MOJAVE 1.5JyQC is shown in Figure 4. All the beaming indicators demonstrate a consistent picture: blazars consistent with neutrino directions have higher Doppler and Lorentz factors, and smaller viewing angles. All of these indicate extremely highly beamed jets, even higher than typical VLBI-selected blazars. A to-scale comparison is shown in Figure 5.
We evaluate the statistical significance of the Doppler factor correlation with neutrino coincidences. Doppler factors are chosen because they are the most direct tracers of the electromagnetic emission beaming. They are also measured more directly from observations, see Homan et al. (2021). We rely on the same statistical testing procedure as in Section 3, namely the shuffling of neutrino event positions in the sky. We choose the Kolmogorov-Smirnov test statistic between the Doppler factor distributions for MOJAVE 1.5JyQC blazars and for neutrino coincidences.
This test is visualized in Figure 4; it demonstrates that the apparent differences are indeed unlikely to arise by chance. The chance probability to obtain the observed correlation is $p=0.3\%$ ($3.0\sigma$) for the P23 event selection, and $p=1\%$ ($2.4\sigma$) for IceCat-1 Gold. We stress that this analysis focuses on the complete VLBI-flux-density limited MOJAVE sample, 1.5JyQC. As in earlier studies, here we chose the cleanest uniform sample of sources for the primary analysis to remain conservative, to keep statistical purity and robustness. Simply extending it to the full MOJAVE dataset would increase the significance level to $3.5\sigma$, see Appendix A. Unlike the 1.5JyQC sample, the full MOJAVE dataset contains sources selected by different criteria accumulated since the start of the program in 1994, see Section 2.1.
The most extreme viewing angle estimates from the MOJAVE observations can be less than $0.1\degree$. There are only three sources with such viewing angles in the full sample (Homan et al., 2021), and one of them is coincident with an IceCube neutrino: 1749+096 with $\theta=0.07\degree$, see Table 1 and Figure 3. We believe that such blazars do indeed have jets pointing almost directly towards us, but recommend treating specific values of these extreme viewing angle estimates with caution. It is possible that some limitations of the generic approach utilized in Homan et al. (2021) manifest in such extreme cases. Additional or alternative ways to infer the jet geometry and the beaming patterns would potentially be useful for more detailed quantitative studies in the future. For example, a very large apparent jet opening angle (Pushkarev et al., 2017) might be used as another independent indicator of a very small viewing angle (e.g., the case of 1424+240, Kovalev et al., in prep.).
Blazars with activity patterns wildly varying over time may also hit limitations of the uniform MOJAVE analysis. In particular, estimates for TXS 0506+056 presented in Table 1 are dominated by its behavior before the major flare in 2018-2022 that followed the neutrino detection. Before that, it was not bright enough to be included into the 1.5JyQC sample, and the lack of activity resulted in low Doppler factor and large viewing angle estimates. Such cases may deserve a dedicated study that takes the flaring behavior into account.
4.2 Beamed neutrino emission implications
Figure 5: A visual comparison of the typical (median) blazar geometry for neutrino-coincident blazars (Table 1) and for the complete MOJAVE 1.5JyQC sample (Homan et al., 2021). The angles are shown to scale with gray lines indicating 1 and 5 degree steps for guidance. The visual scaling of the velocity and beaming factors is arbitrary.
Findings presented in Section 3 and Section 4.1 provide an evidence for a high beaming of the jet radio emission in neutrino-associated blazars at a range of scales. From a fraction of a parsec (VLBI core brightness, Doppler factors) to tens of parsecs (total VLBI flux density, kinematics measurements), we see that blazars coincident with neutrino events are even more extreme than other VLBI blazars (Figure 5). This unambiguously implies that the neutrino emission itself is highly anisotropic, these particles are predominantly emitted along the jet axis. They are detected on Earth preferentially from sources with the jet pointing straight towards us. The beaming appears to affect both neutrinos and photons in a similar way, but detailed quantification of this effect requires a more reliable identification of individual neutrino-emitting blazars and their properties.
[…]
Neutrino production at $\sim$ parsec distances from the central black hole is consistent with the likely scenario of $p\gamma$ interactions within the jet. There, the required X-ray target photon field can be produced either by the jet itself (self- or external Compton, Plavin et al., 2021; Kalashev et al., 2023) or by other regions of the AGN, such as the corona (e.g., Inoue et al., 2019; Murase et al., 2020; Sotirov, 2024). The exact origin of these photons remains an open question, together with the specific mechanism accelerating protons to PeV energies.
We note that our findings do not preclude the presence of other, non-beamed, sources of neutrinos. The neutrino–VLBI flux density and neutrino–Doppler factor correlations imply that there is a population of blazars that emit neutrinos, and emit them predominantly along the jet axis. Non-beamed and non-jet-related neutrino production, such as possible for Seyfert-type AGNs like NGC1068 (IceCube Collaboration et al., 2022; Abbasi et al., 2024b; Neronov et al., 2024), may form another population of sources leading to a diverse extragalactic neutrino sky.
Heavily flaring blazars producing neutrino during their flares may potentially form another population of neutrino sources that our analysis is not sensitive to. Earlier studies reported observational evidence for neutrinos coming preferentially during electromagnetic flux increases, especially in the radio band: P20; Allakhverdyan et al. (2024b); Hovatta et al. (2021); Albert et al. (2024); Kouch et al. (2024). Those sources may or may not demonstrate extreme jet beaming and persistently bright radio emission. For example, TXS 0506+056 exhibited barely any activity before the neutrino arrival and the following major radio flare in 2018 – 2022. This likely affects its beaming estimates, see Section 4.1. Meanwhile, PKS 1502+106 is another heavily flaring blazar that was originally selected for its highest temporal neutrino-radio correlation in P23 and later studied as a likely neutrino counterpart (e.g. Oikonomou et al., 2021; Britzen et al., 2021). Unlike TXS 0506+056, this blazar demonstrates an extremely strong beaming in our analysis, see Table 1.
5 Conclusions
In this work, we studied direct observational tracers of jet beaming in blazars coincident with high-energy neutrinos detected with the IceCube. The blazars located around neutrino arrival directions were found to exhibit extremely high beaming with median Doppler factor of $D\approx 30$ and a median viewing angle of $\theta\approx 0.7\degree$. These blazars are even more extreme than already highly beamed VLBI-selected ones that typically exhibit $D\approx 15$, $\theta\approx 3\degree$ (Figure 5). The statistical significance of the Doppler factor versus neutrino correlation is $3\sigma$ and $2.4\sigma$ for different IceCube event selections (see Section 4 and Figure 4).
This population study provides yet another piece of evidence that bright blazars are strong neutrino emitters, and the observed neutrino-radio blazar correlation represents a genuine astrophysical connection, not a statistical fluctuation. The effect is consistent across different IceCube event selections and manifests itself in all observational indicators of the jet beaming as shown in Figure 2 and Figure 4. In agreement with earlier studies, enlarging the reported IceCube directional uncertainties (by a previously determined value of $0.78\degree$) is necessary to unravel this connection, see Section 3.
The observational connection between jet electromagnetic emission beaming and detected neutrinos requires the neutrino emission to be highly beamed as well. The beaming effects appear to be comparable for electromagnetic and neutrino emission, but quantifying them would require both more detailed modeling (to handle observational biases and the intrinsic diversity of blazars) and robust individual neutrino associations. Neutrino beaming implies that the multi-PeV protons producing them exhibit a strong bulk motion along the jet direction. In typical jet acceleration profiles, this motion is only attained at parsec scales, potentially providing a lower limit on the distance from the central black hole to the neutrino production region.
Evaluating direct jet beaming indicators of neutrino-coincident blazars, such as the Doppler factor, already draws a consistent picture of their jets as extremely relativistic and aligned with the line of sight. These indicators are challenging to measure and can only be determined for a limited subset of blazars with great observational coverage. Directly observable properties such as the parsec-scale emission strength measured by VLBI remain well-suited for larger scale searches for neutrino – blazar associations.
Acknowledgements
We thank Eduardo Ros for comments on the manuscript, Markus Böttcher, Anna Franckowiak, and Emma Kun for productive discussions. A.V.P. is a postdoctoral fellow at the Black Hole Initiative, which is funded by grants from the John Templeton Foundation (grants 60477, 61479, 62286) and the Gordon and Betty Moore Foundation (grant GBMF-8273). Y.Y.K. is supported by the MuSES project which has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 Research and Innovation Programme (grant agreement No 101142396). The work of S.V.T. is supported in the framework of the State project “Science” by the Ministry of Science and Higher Education of the Russian Federation under the contract 075-15-2024-541. The opinions expressed in this work are those of the authors and do not necessarily reflect the views of these Foundations.
This research has made use of the NASA/IPAC Extragalactic Database, which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology. This research has made use of data from the Radio Fundamental Catalog (Petrov & Kovalev, 2025, doi:10.25966/dhrk-zh08). This research made use of the data from the MOJAVE database maintained by the MOJAVE team (Lister et al., 2018).
Appendix A Beaming of neutrino-coincident blazars in the full MOJAVE sample
In the main analysis of this paper, Section 4.1, we utilize the uniformly-selected subsample of the full set of blazars observed within the MOJAVE program (1.5JyQC). That subset is purely flux-density limited (Section 2.1), leading to the most robust findings. Still, it is potentially instructive to look at the full MOJAVE blazar sample consisting of 447 sources with measured Doppler boosting (Homan et al., 2021). The comparison is shown in Figure A1. The presence of relatively fainter blazars makes the neutrino-coincident ones stand out even more than in the 1.5JyQC analysis (Figure 4). This effect highlights the importance of population studies, large-sample survey and monitoring programs.
Figure A1: Comparison of observational jet beaming indicators (Doppler $D$ and Lorentz $\Gamma$ factors, viewing angle $\theta$) for neutrino-coincident blazars and for the full sample of blazars observed in MOJAVE. For comparison, see Figure 4 that only shows 1.5JyQC sources.

### v2

@@
 Blazars were the first high-energy (TeVs-PeVs) astrophysical neutrino sources to be discovered (IceCube Collaboration et al., 2018).
-Later, neutrinos have also been detected from other sources, including the Milky Way Galaxy (Kovalev et al., 2022; Icecube Collaboration et al., 2023; Albert et al., 2023; Allakhverdyan et al., 2024a), nearby Seyfert-type Active Galactic Nuclei (AGNs, IceCube Collaboration et al., 2022; Neronov et al., 2024), and Tidal Disruption Events in AGNs (TDEs, van Velzen et al., 2024).
+Later, neutrinos have also been detected from other sources, for instance the Milky Way Galaxy (Y.
+Y.
+Kovalev et al., 2022; Icecube Collaboration et al., 2023; A.
+Albert et al., 2023; V.
+A.
+Allakhverdyan et al., 2024a) and nearby Seyfert-type Active Galactic Nuclei (AGNs, R.
+Abbasi et al., 2022; A.
+Neronov et al., 2024).
 Still, blazars remain a very prominent class of bright neutrino emitters.
-There are more and more population-based studies resulting in neutrinos associated with AGNs selected by their compact radio flux (basically, blazars; e.g., P23; Kouch et al.
-2024; Albert et al.
-2024; Suray & Troitsky 2024), X-ray flux (Kun et al.
-2024; Khatee Zathul et al.
-2024; Neronov et al.
-2024; blazars in Plavin et al.
-2024), or using more specialized blazar selection methods (e.g., Kadler et al., 2016; Giommi et al., 2020; Buson et al., 2022).
+There are more and more population-based studies resulting in neutrinos associated with AGNs selected by their compact radio flux (basically, blazars; e.g., A.
+V.
+Plavin et al.
+2023; P.
+M.
+Kouch et al.
+2024; A.
+Albert et al.
+2024; A.
+Suray & S.
+Troitsky 2024), X-ray flux (E.
+Kun et al.
+2024; A.
+Neronov et al.
+2024; blazars in A.
+V.
+Plavin et al.
+2024), or using more specialized blazar selection methods (e.g., M.
+Kadler et al., 2016; P.
+Giommi et al., 2020; S.
+Buson et al., 2022).
+In this work, we adopt the commonly used definition of “blazar”, referring to an AGN with a powerful jet aligned close to our line of sight.
 Recent studies (e.g.
-P23; Kouch et al.
+A.
+V.
+Plavin et al.
+2023; P.
+M.
+Kouch et al.
 2024) have specifically demonstrated a strong observational connection of radio-bright blazars being also bright neutrino sources.
@@
 Those results already provide an indication that neutrinos are emitted predominantly along the jet axis in blazars; more direct studies should help quantifying this effect.
-Furthermore, note that the strong neutrino-blazars association is not found in every analysis, depending on the handling of directional uncertainties (see discussion in Abbasi et al., 2023; Kouch et al., 2024).
+Furthermore, note that the strong neutrino-blazars association is not found in every analysis, depending on the handling of directional uncertainties (see discussion in R.
+Abbasi et al., 2023; P.
+M.
+Kouch et al., 2024).
 Utilizing additional observables is thus crucial to refine and potentially confirm the connection.
@@
 There are 83 blazars from the RFC catalog coincident with at least one IceCube neutrino; note that they include both real physical associations and chance coincidences, see Figure 2.
-Neutrino-coincident MOJAVE 1.5JyQC (Section 2.1) blazars are labeled here; see Figure 3 for their VLBI images and Table 1 for their properties.
+Neutrino-coincident MOJAVE 1.5JyQC (Section 2.1) blazars are labeled here; see Figure A1 for their VLBI images and Table 1 for their properties.
 2.1 Radio observations of parsec-scale emission
@@
 For Doppler boosting and jet viewing angle measurements, we rely on the results of the MOJAVE (Monitoring Of Jets in Active galactic nuclei with VLBA Experiments) team.
-These parameters cannot be measured directly for blazars; the MOJAVE team has suggested (Homan et al., 2006) and applied (Homan et al., 2021) a procedure to infer them from VLBI monitoring data; we use those measurements in this paper.
-The full MOJAVE dataset is assembled with a range of selection criteria: bright VLBI-selected blazars, compact symmetric objects, Fermi-selected blazars, low-luminosity AGNs (Lister et al., 2011, 2019, 2020, 2013).
-Similarly to our earlier studies, we focus on a uniform flux-density limited subsample of the MOJAVE program sources in the analysis: the 1.5 Jy Quarter Century Sample (1.5JyQC; defined in Lister et al.
+These parameters cannot be measured directly for blazars; the MOJAVE team has suggested (D.
+C.
+Homan et al., 2006) and applied (D.
+C.
+Homan et al., 2021) a procedure to infer them from VLBI monitoring data; we use those measurements in this paper.
+The full MOJAVE dataset is assembled with a range of selection criteria: bright VLBI-selected blazars, compact symmetric objects, Fermi-selected blazars, low-luminosity AGNs (M.
+L.
+Lister et al., 2011, 2019, 2020, 2013).
+Similarly to our earlier studies, we focus on a uniform flux-density limited subsample of the MOJAVE program sources in the analysis: the 1.5 Jy Quarter Century Sample (1.5JyQC; defined in M.
+L.
+Lister et al.
 2019).
@@
 The first catalog is the sample of high-energy alert-like neutrino detections we collected before, most recently in P23.
-Earlier, the same selection criteria were utilized in P20 and Hovatta et al.
+Earlier, the same selection criteria were utilized in P20 and T.
+Hovatta et al.
 (2021).
@@
 The estimated signalness is $\geq 50\%$ for all these events; their neutrino energy estimates range from 56 TeV to 6 PeV.
-For the analyses performed in this work, we enlarge the original IceCube directional 90% uncertainties by $0.78\degree$ isotropically in all directions.
-This enlargement was determined in P23 to be optimal for association searches, and we keep the same value to avoid multiple trials in the statistical analysis.
-A similar enlargement was found in Kouch et al.
+The sky localization of IceCube events from both samples is shown in Figure 1.
+Blazars coincident with IceCube events are highlighted in the figure, and those from the MOJAVE 1.5JyQC sample are labeled.
+See Table 1 for the properties of neutrino-coincident blazars from the full MOJAVE sample.
+Note how the change of IceCube event parameters between different reconstructions leads to slightly different lists of events and of their blazar coincidences.
+Evaluating the reconstruction differences in details is out of scope of this paper; here, we use these two catalogs as an extra robustness check of our results, reporting only conclusions consistent between the two.
+For the analyses performed in this work, we enlarge the original 90% directional uncertainties reported in IceCube catalogs by $0.78\degree$ isotropically in all directions.
+This enlargement magnitude was determined in P23 to be optimal for association searches, and we keep the same value to avoid multiple trials in the statistical analysis.
+As a guidance, Table 1 marks the three coincidences that don’t require this enlargement.
+A similar enlargement was found in P.
+M.
+Kouch et al.
 (2024); see that paper, P23, and Section 3 for a more detailed discussion.
-The sky localization of IceCube events from both samples is shown in Figure 1.
-Blazars coincident with IceCube events are highlighted in the figure, and those from the MOJAVE 1.5JyQC sample are labeled; see Table 1 for their properties.
-Figure 2: The distribution of VLBI 8 GHz flux density for blazars coincident with IceCube neutrinos, shown as histograms and cumulative distribution plots.
-Similarly to Figure 1, all VLBI blazars from the RFC catalog are shown in black exactly the same in the left and right panels.
-Neutrino coincidences are shown in red, using the P23 events sample on the left and the IceCat-1 on the right (see Section 2.2).
 3 Neutrino-blazar correlation
@@
 (1) (2) (3) (4) (5) (6) (7) (8) (9) (10) (11)
-2015--09--26 IC150926A J1256--0547 1253--055 $0.536$ $15$ $19$ yes $140$ $72$ $0.12$
-2022--03--03 J1751$+$0939 1749$+$096 $0.322$ $2.8$ $4$ yes $100$ $51$ $0.07$
-2019--07--30 IC190730A J1504$+$1029 1502$+$106 $1.838$ $1.5$ $1.4$ yes $65$ $35$ $0.44$
-2016--01--28 IC160128A J1733--1304 1730--130 $0.902$ $4$ $4.5$ yes $48$ $32$ $1$
-2015--08--12 IC150812B J2148$+$0657 2145$+$067 $0.999$ $6.6$ $7.5$ yes $31$ $16$ $0.36$
-2015--08--31 IC150831A J0336$+$3218 0333$+$321 $1.259$ $1.6$ $1.7$ yes $30$ $18$ $1.4$
-2012--05--15 IC120515A J1310$+$3220 1308$+$326 $0.997$ $1.9$ $2.1$ yes $24$ $28$ $2.3$
-2022--02--05 IC220205B J1743--0350 1741--038 $1.054$ $4$ $5.3$ yes $22$
-2011--09--30 J1743--0350 1741--038 $1.054$ $4$ $5.3$ yes $22$
-2013--10--14 IC131014A J0211$+$1051 0208$+$106 $0.2$ $0.66$ $1.1$ yes $6.3$ $5.1$ $8.8$
-2020--11--30 IC201130A J0206--1150 0203--120 $1.663$ $0.28$ $0.3$ no $35$ $20$ $1$
-2017--09--22 IC170922A J0505$+$0459 0502$+$049 $0.954$ $0.7$ $0.93$ no $25$ $14$ $1.6$
-IC221223A J2311$+$3425 2308$+$341 $1.817$ $0.69$ $0.86$ no $23$
-IC201221A J1724$+$4004 1722$+$401 $1.049$ $0.5$ $0.58$ no $9.8$ $61$ $3.2$
-2010--10--09 J2200$+$1030 2157$+$102 $0.172$ $0.19$ $0.22$ no $5$
-2017--09--22 IC170922A J0509$+$0541 0506$+$056 $0.3365$ $0.42$ $0.46$ no $1.8$ $1.5$ $33$
-2011--07--14 IC110714A J0432$+$4138 0429$+$415 $1.022$ $1.4$ $1$ no $0.13$ $54$ $31$
+2015-09-26b IC150926Ab J1256-0547 1253-055 $0.536$ $15$ 1919 yes $140$ $72$ $0.12$
+2022-03-03b J1751++0939 1749++096 $0.322$ $2.8$ 44 yes $100$ $51$ $0.07$
+2019-07-30a IC190730Aa J1504++1029 1502++106 $1.838$ $1.5$ $1.4$ yes $65$ $35$ $0.44$
+2016-01-28b IC160128Ab J1733-1304 1730-130 $0.902$ 44 $4.5$ yes $48$ $32$ 11
+2015-08-12b IC150812Bb J2148++0657 2145++067 $0.999$ $6.6$ $7.5$ yes $31$ $16$ $0.36$
+2015-08-31b IC150831Ab J0336++3218 0333++321 $1.259$ $1.6$ $1.7$ yes $30$ $18$ $1.4$
+2012-05-15b IC120515Ab J1310++3220 1308++326 $0.997$ $1.9$ $2.1$ yes $24$ $28$ $2.3$
+2022-02-05b IC220205Bb J1743-0350 1741-038 $1.054$ 44 $5.3$ yes $22$
+2011-09-30b J1743-0350 1741-038 $1.054$ 44 $5.3$ yes $22$
+2013-10-14b IC131014Ab J0211++1051 0208++106 $0.2$ $0.66$ $1.1$ yes $6.3$ $5.1$ $8.8$
+2020-11-30a IC201130Aa J0206-1150 0203-120 $1.663$ $0.28$ $0.3$ no $35$ 2020 11
+2017-09-22b IC170922Ab J0505++0459 0502++049 $0.954$ $0.7$ $0.93$ no $25$ $14$ $1.6$
+IC221223Ab J2311++3425 2308++341 $1.817$ $0.69$ $0.86$ no $23$
+IC201221Ab J1724++4004 1722++401 $1.049$ $0.5$ $0.58$ no $9.8$ $61$ $3.2$
+2010-10-09b J2200++1030 2157++102 $0.172$ $0.19$ $0.22$ no 55
+2017-09-22a IC170922Aa J0509++0541 0506++056 $0.3365$ $0.42$ $0.46$ no $1.8$ $1.5$ $33$
+2011-07-14b IC110714Ab J0432++4138 0429++415 $1.022$ $1.4$ 11 no $0.13$ $54$ $31$
 Note.
-— Columns are as follows: (1) — IceCube neutrino arrival date for events included in the P23 sample; (2) — IceCat event ID for events included in IceCat-1; see Section 2.2 for a definition of these samples; (3), (4) — coordinate-based blazar names, with the J2000 name taken directly from the RFC (Petrov & Kovalev, 2025); (5) — redshift $z$ extracted from the NASA/IPAC Extragalactic Database (NED); (6), (7) — parsec-scale radio flux densities measured by VLBI, as presented in the RFC catalog; their typical uncertainty is about $10\%$, dominated by the amplitude calibration; (8) — flag indicating whether the blazar is included in the uniformly-selected MOJAVE 1.5JyQC sample (Section 2.1); those sources are also labeled in Figure 1, and their VLBI images are shown in Figure 3; (9), (10), (11) — beaming indicators, Doppler $D$ and Lorentz $\Gamma$ factors and the viewing angle $\theta$, as measured within the MOJAVE program (Homan et al., 2021); see Section 4.1 for a discussion of assumptions and uncertainties of these measurements.
+— Columns are as follows: (1) — IceCube neutrino arrival date for events included in the P23 sample; (2) — IceCat event ID for events included in IceCat-1; see Section 2.2 for a definition of these samples; a indicates neutrino-blazar matches within original reported uncertainties, and b those matches that require extending original uncertainties, following Section 3.
+(3), (4) — coordinate-based blazar names, with the J2000 name taken directly from the RFC (L.
+Y.
+Petrov & Y.
+Y.
+Kovalev, 2025); (5) — redshift $z$ extracted from the NASA/IPAC Extragalactic Database (NED); (6), (7) — parsec-scale radio flux densities measured by VLBI, as presented in the RFC catalog; their typical uncertainty is about $10\%$, dominated by the amplitude calibration; (8) — flag indicating whether the blazar is included in the uniformly-selected MOJAVE 1.5JyQC sample (Section 2.1); those sources are also labeled in Figure 1, and their VLBI images are shown in Figure A1; (9), (10), (11) — beaming indicators, Doppler $D$ and Lorentz $\Gamma$ factors and the viewing angle $\theta$, as measured within the MOJAVE program (D.
+C.
+Homan et al., 2021); see Section 4.1 for a discussion of assumptions and uncertainties of these measurements.
 The blazars are ordered by the inclusion in MOJAVE 1.5JyQC (8) and by the Doppler factor measurement (9).
@@
 After having chosen the test statistic, we calculate the probability that its observed value could have arisen by chance, assuming no directional neutrino-blazar correlation.
-For that, we randomize the Right Ascensions of all IceCube events $10^{6}$ times, motivated by the instrument sensitivity being dependent on the Declination only (used in Aartsen et al., 2017, and elsewhere).
+For that, we randomize the Right Ascensions of all IceCube events $10^{6}$ times, motivated by the instrument sensitivity being dependent on the Declination only (used in M.
+G.
+Aartsen et al., 2017, and elsewhere).
 Then, the chance probability is the fraction of random realizations where the average VLBI flux of matching blazars is higher than the observed value.
 There are no free parameters to be optimized in this procedure, and we can immediately interpret the resulting fraction as the $p$-value.
+Figure 2: The distribution of VLBI 8 GHz flux density for blazars coincident with IceCube neutrinos, shown as histograms and cumulative distribution plots.
+Similarly to Figure 1, all VLBI blazars from the RFC catalog are shown in black exactly the same in the left and right panels.
+Neutrino coincidences are shown in red, using the P23 events sample on the left and the IceCat-1 on the right (see Section 2.2).
 The P23 IceCube event selection.
@@
 We find such an agreement between results obtained from different event reconstructions, selections, and categorization approaches quite remarkable.
-Note that updated or more detailed event reconstructions are not necessarily better in every aspect: see demonstrations and discussion in the review Troitsky (2024) and in Sommani et al.
+Note that updated or more detailed event reconstructions are not necessarily better in every aspect: see demonstrations and discussion in the review S.
+Troitsky (2024) and in G.
+Sommani et al.
 (2024).
-For example, a reconstruction update lead to a decrease of the TXS 0506+056 flare significance (IceCube Collaboration et al., 2021; Abbasi et al., 2023).
+For example, a reconstruction update lead to a decrease of the TXS 0506+056 flare significance (IceCube Collaboration et al., 2021; R.
+Abbasi et al., 2023).
 A detailed analysis and interpretation of the differences between various IceCube event reconstructions is out of scope of this work.
@@
 Cross-check without enlarging IceCube errors.
-As detailed in Section 2.2, we enlarge original IceCube event uncertainties by $0.78\degree$ for all calculations in this paper, following the approach of P23 and Kouch et al.
+As detailed in Section 2.2, we enlarge original IceCube event uncertainties by $0.78\degree$ for all calculations in this paper, following the approach of P23 and P.
+M.
+Kouch et al.
 (2024) and avoiding multiple comparisons issues.
@@
 The spatial correlation shown above is heavily driven by blazars that lie slightly outside the original reported 90% directional errors.
-This is still consistent with the observed correlation caused by a physical connection.
+This is still consistent with the observed correlation caused by a physical connection, see P.
+M.
+Kouch et al.
+(2025) for a sensitivity and robustness analysis.
 Such an effect could result from unaccounted systematic uncertainties or other discrepancies between estimated and actual error regions.
-A detailed dedicated study can only be performed by the IceCube collaboration, utilizing all the available data about each detection; see Lagunas Gualda et al.
+A detailed dedicated study can only be performed by the IceCube collaboration, utilizing all the available data about each detection; see C.
+Lagunas Gualda et al.
 (2021) for a discussion and an overview of challenges.
 Further in the future, it will be crucial to compare the brightest neutrino sources in the sky between different neutrino observatories.
-The cross-check we performed here does not replace the detailed analysis performed in P23 and Kouch et al.
-(2024), but remains consistent with those works and results reported by the IceCube (Abbasi et al., 2023, 2024a).
+The cross-check we performed here does not replace the detailed analysis performed in P23 and P.
+M.
+Kouch et al.
+(2024), but remains consistent with those works and results reported by R.
+Abbasi et al.
+(2023, 2024a); C.
+Bellenghi et al.
+(2023); J.
+Kuhlmann & F.
+Capel (2025).
 4 Geometry and kinematics of neutrino-associated blazars
-Figure 3: MOJAVE VLBI images for the nine 1.5JyQC blazars coincident with IceCube neutrinos from both samples of Section 2.
-We show here averages of multi-decade individual VLBI images, which provide a better dynamic range to trace the overall structure of the jet.
-These stacked images are available from the MOJAVE website (https://www.cv.nrao.edu/MOJAVE/), see details in Pushkarev et al.
-(2017).
-For each blazar, we show the redshift $z$, the Doppler factor $D$ and the viewing angle $\theta$ estimates.
-Blazars are ordered by their Doppler factor, same as in Table 1.
 Figure 3: Comparison of observational jet beaming indicators (Doppler $D$ and Lorentz $\Gamma$ factors, viewing angle $\theta$) for neutrino-coincident blazars and for the complete MOJAVE 1.5JyQC sample (see Section 2 and Table 1).
@@
 The association of high-energy neutrinos with VLBI-bright blazars (Section 3) already suggests that beaming plays a major role in neutrino production and detection: neutrinos are predominantly emitted along the direction of the jet.
-Indeed, VLBI-selected AGNs are dominated by galaxies with jets pointing towards us, their typical viewing angles are $\theta=2\dots 3\degree$ (Lister et al., 2019; Homan et al., 2021).
-If neutrino emission were isotropic, we would expect to see thousands ($30004\pi/\pi\theta^{2}~{}\sim~{}$ 3000) of similar AGNs pointed in other directions for each neutrino-VLBI association.
+Indeed, VLBI-selected AGNs are dominated by galaxies with jets pointing towards us, their typical viewing angles are $\theta=2\dots 3\degree$ (M.
+L.
+Lister et al., 2019; D.
+C.
+Homan et al., 2021).
+If neutrino emission were isotropic, we would expect to see thousands ($4\pi/\pi\theta^{2}~\sim~3000$) of similar AGNs pointed in other directions for each neutrino-VLBI association.
 This is clearly incompatible with the observed number of high-energy neutrinos.
@@
 $\begin{split}\Gamma&=\frac{1}{\sqrt{1-\beta^{2}}}\\ D&=\frac{1}{\Gamma(1-\beta\cos\theta)}\end{split}$ (1)
-They are measured from VLBI observations within the MOJAVE program (Section 2.1, Homan et al.
+They are measured from VLBI observations within the MOJAVE program (Section 2.1, D.
+C.
+Homan et al.
 2021).
@@
 It can be challenging to calculate precise uncertainties of these measurements.
-Still, as a cross-check, Doppler factors resulting from MOJAVE measurements were found to be consistent with independent estimates obtained via different techniques, such as radio variability (Liodakis et al., 2018; Homan et al., 2021).
+Still, as a cross-check, Doppler factors resulting from MOJAVE measurements were found to be consistent with independent estimates obtained via different techniques, such as radio variability (I.
+Liodakis et al., 2018; D.
+C.
+Homan et al., 2021).
 We probe the jet beaming — neutrino emission connection by comparing these beaming indicators (Doppler factor, plasma velocity, viewing angle) between neutrino-coincident blazars and the overall sample.
@@
 Doppler factors are chosen because they are the most direct tracers of the electromagnetic emission beaming.
-They are also measured more directly from observations, see Homan et al.
+They are also measured more directly from observations, see D.
+C.
+Homan et al.
 (2021).
@@
 The most extreme viewing angle estimates from the MOJAVE observations can be less than $0.1\degree$.
-There are only three sources with such viewing angles in the full sample (Homan et al., 2021), and one of them is coincident with an IceCube neutrino: 1749+096 with $\theta=0.07\degree$, see Table 1 and Figure 3.
+There are only three sources with such viewing angles in the full sample (D.
+C.
+Homan et al., 2021), and one of them is coincident with an IceCube neutrino: 1749+096 with $\theta=0.07\degree$, see Table 1 and Figure A1.
 We believe that such blazars do indeed have jets pointing almost directly towards us, but recommend treating specific values of these extreme viewing angle estimates with caution.
-It is possible that some limitations of the generic approach utilized in Homan et al.
+It is possible that some limitations of the generic approach utilized in D.
+C.
+Homan et al.
 (2021) manifest in such extreme cases.
 Additional or alternative ways to infer the jet geometry and the beaming patterns would potentially be useful for more detailed quantitative studies in the future.
-For example, a very large apparent jet opening angle (Pushkarev et al., 2017) might be used as another independent indicator of a very small viewing angle (e.g., the case of 1424+240, Kovalev et al., in prep.).
+For example, a very large apparent jet opening angle (A.
+B.
+Pushkarev et al., 2017) might be used as another independent indicator of a very small viewing angle (e.g., the case of PKS 1424+240, Y.
+Y.
+Kovalev et al.
+2025).
 Blazars with activity patterns wildly varying over time may also hit limitations of the uniform MOJAVE analysis.
@@
 4.2 Beamed neutrino emission implications
-Figure 5: A visual comparison of the typical (median) blazar geometry for neutrino-coincident blazars (Table 1) and for the complete MOJAVE 1.5JyQC sample (Homan et al., 2021).
-The angles are shown to scale with gray lines indicating 1 and 5 degree steps for guidance.
-The visual scaling of the velocity and beaming factors is arbitrary.
+Figure 4: Schematic illustration comparing the typical geometries for blazars in the MOJAVE 1.5JyQC sample (top; D.
+C.
+Homan et al.
+2021) and neutrino-coincident blazars (bottom; Table 1).
+Angles are shown to scale, with gray lines indicating $1\degree$ and $5\degree$ increments for guidance.
+All quantities depicted are medians within their respective samples.
+The visual scaling of velocity and beaming factors is arbitrary.
 Findings presented in Section 3 and Section 4.1 provide an evidence for a high beaming of the jet radio emission in neutrino-associated blazars at a range of scales.
@@
 Neutrino production at $\sim$ parsec distances from the central black hole is consistent with the likely scenario of $p\gamma$ interactions within the jet.
-There, the required X-ray target photon field can be produced either by the jet itself (self- or external Compton, Plavin et al., 2021; Kalashev et al., 2023) or by other regions of the AGN, such as the corona (e.g., Inoue et al., 2019; Murase et al., 2020; Sotirov, 2024).
+There, the required X-ray target photon field can be produced either by the jet itself (self- or external Compton, A.
+V.
+Plavin et al., 2021; O.
+Kalashev et al., 2023; D.
+F.
+G.
+Fiorillo et al., 2025) or by other regions of the AGN, such as the corona (e.g., Y.
+Inoue et al., 2019; K.
+Murase et al., 2020; S.
+Sotirov, 2024).
 The exact origin of these photons remains an open question, together with the specific mechanism accelerating protons to PeV energies.
@@
 The neutrino–VLBI flux density and neutrino–Doppler factor correlations imply that there is a population of blazars that emit neutrinos, and emit them predominantly along the jet axis.
-Non-beamed and non-jet-related neutrino production, such as possible for Seyfert-type AGNs like NGC1068 (IceCube Collaboration et al., 2022; Abbasi et al., 2024b; Neronov et al., 2024), may form another population of sources leading to a diverse extragalactic neutrino sky.
+Non-beamed and non-jet-related neutrino production, such as possible for Seyfert-type AGNs like NGC1068 (R.
+Abbasi et al., 2022, 2024b; A.
+Neronov et al., 2024), may form another population of sources leading to a diverse extragalactic neutrino sky.
 Heavily flaring blazars producing neutrino during their flares may potentially form another population of neutrino sources that our analysis is not sensitive to.
-Earlier studies reported observational evidence for neutrinos coming preferentially during electromagnetic flux increases, especially in the radio band: P20; Allakhverdyan et al.
-(2024b); Hovatta et al.
-(2021); Albert et al.
-(2024); Kouch et al.
+Earlier studies reported observational evidence for neutrinos coming preferentially during electromagnetic flux increases, especially in the radio band: P20; V.
+A.
+Allakhverdyan et al.
+(2024b); T.
+Hovatta et al.
+(2021); A.
+Albert et al.
+(2024); P.
+M.
+Kouch et al.
 (2024).
@@
 Meanwhile, PKS 1502+106 is another heavily flaring blazar that was originally selected for its highest temporal neutrino-radio correlation in P23 and later studied as a likely neutrino counterpart (e.g.
-Oikonomou et al., 2021; Britzen et al., 2021).
+F.
+Oikonomou et al., 2021; S.
+Britzen et al., 2021).
 Unlike TXS 0506+056, this blazar demonstrates an extremely strong beaming in our analysis, see Table 1.
-5 Conclusions
+5 Summary
 In this work, we studied direct observational tracers of jet beaming in blazars coincident with high-energy neutrinos detected with the IceCube.
@@
 In agreement with earlier studies, enlarging the reported IceCube directional uncertainties (by a previously determined value of $0.78\degree$) is necessary to unravel this connection, see Section 3.
-The observational connection between jet electromagnetic emission beaming and detected neutrinos requires the neutrino emission to be highly beamed as well.
+The observational connection between the jet electromagnetic emission beaming and detected neutrinos requires the neutrino emission to be highly beamed as well.
 The beaming effects appear to be comparable for electromagnetic and neutrino emission, but quantifying them would require both more detailed modeling (to handle observational biases and the intrinsic diversity of blazars) and robust individual neutrino associations.
@@
 Evaluating direct jet beaming indicators of neutrino-coincident blazars, such as the Doppler factor, already draws a consistent picture of their jets as extremely relativistic and aligned with the line of sight.
-These indicators are challenging to measure and can only be determined for a limited subset of blazars with great observational coverage.
+This effect is demonstrated here by population analysis, and complemented by dedicated studies of individual sources.
+In particular, see Y.
+Y.
+Kovalev et al.
+(2025) that finds and explores the extreme jet beaming in the neutrino-associated TeV blazar with low VLBI apparent speed PKS 1424+240 (R.
+Abbasi et al., 2022).
+Note that these beaming indicators are challenging to measure and can only be determined for a limited subset of blazars with great observational coverage.
 Directly observable properties such as the parsec-scale emission strength measured by VLBI remain well-suited for larger scale searches for neutrino – blazar associations.
 Acknowledgements
-We thank Eduardo Ros for comments on the manuscript, Markus Böttcher, Anna Franckowiak, and Emma Kun for productive discussions.
+We thank Eduardo Ros and an anonymous referee for comments on the manuscript, Markus Böttcher, Anna Franckowiak, and Emma Kun for productive discussions.
 A.V.P.
@@
 Y.Y.K.
-is supported by the MuSES project which has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 Research and Innovation Programme (grant agreement No 101142396).
+was supported by the MuSES project, which has received funding from the European Union (ERC grant agreement No 101142396).
 The work of S.V.T.
 is supported in the framework of the State project “Science” by the Ministry of Science and Higher Education of the Russian Federation under the contract 075-15-2024-541.
-The opinions expressed in this work are those of the authors and do not necessarily reflect the views of these Foundations.
+The views and opinions expressed in this work are those of the authors and do not necessarily reflect the views of these Foundations.
 This research has made use of the NASA/IPAC Extragalactic Database, which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology.
-This research has made use of data from the Radio Fundamental Catalog (Petrov & Kovalev, 2025, doi:10.25966/dhrk-zh08).
-This research made use of the data from the MOJAVE database maintained by the MOJAVE team (Lister et al., 2018).
+This research has made use of data from the Radio Fundamental Catalog (L.
+Y.
+Petrov & Y.
+Y.
+Kovalev, 2025, doi:10.25966/dhrk-zh08).
+This research made use of the data from the MOJAVE database maintained by the MOJAVE team (M.
+L.
+Lister et al., 2018).
+Code to reproduce most of the figures from this manuscript can be found at https://github.com/aplavin/astroplots.jl/tree/master/nubeam25.
 Appendix A Beaming of neutrino-coincident blazars in the full MOJAVE sample
+Figure A1: MOJAVE VLBI images for the nine 1.5JyQC blazars coincident with IceCube neutrinos from both samples of Section 2.
+We show here averages of multi-decade individual VLBI images, which provide a better dynamic range to trace the overall structure of the jet.
+These stacked images are available from the MOJAVE website (https://www.cv.nrao.edu/MOJAVE/), see details in A.
+B.
+Pushkarev et al.
+(2017).
+For each blazar, we show the redshift $z$, the Doppler factor $D$ and the viewing angle $\theta$ estimates.
+Blazars are ordered by their Doppler factor, same as in Table 1.
 In the main analysis of this paper, Section 4.1, we utilize the uniformly-selected subsample of the full set of blazars observed within the MOJAVE program (1.5JyQC).
 That subset is purely flux-density limited (Section 2.1), leading to the most robust findings.
-Still, it is potentially instructive to look at the full MOJAVE blazar sample consisting of 447 sources with measured Doppler boosting (Homan et al., 2021).
+The VLBI images of the nine 1.5JyQC blazars coincident with neutrinos are shown in Figure A1.
+Still, it is potentially instructive to look at the full MOJAVE blazar sample consisting of 447 sources with measured Doppler boosting (D.
+C.
+Homan et al., 2021).
 The comparison is shown in Figure A2.